(-3,4)and (9,12)
Determine the co-ordinates X intercept and y intercept

Answer by mananth(16946)

x1 y1 x2 y2
-3 4 9 12

slope m = (y2-y1)/(x2-x1)
(12-4)/(9-(-3))
(8/12)
m=0.67

Plug value of the slope and point (-3,4) in
Y =mx+b
4.00=-2+ b
b=4-2
b=6
So the equation will be
Y =0.67x +6
when x=0 y= 6 ( 0,6),y intercept point
when y=0,-0.67x=6
x=-6/0.67
x=-8.96 (-8.96,0),x intercept point